A tugboat goes 180 miles upstream in 15 hours. The return trip downstream takes 10 hours. Find the speed of the tugboat wihtout a current and the speed of the current.
Answer by stanbon(57410)

A tugboat goes 180 miles upstream in 15 hours. The return trip downstream takes 10 hours. Find the speed of the tugboat wihtout a current and the speed of the current.
--------------------------
Upstream DATA:
distance = 180 miles ; time = 15 hrs ; rate = 180/15 = 12 mph
-----
Downstream DATA:
distance = 180 miles ; time = 10 hrs ; rate = 180/10 = 18 mph
------
Equation:
Upstream: b - c = 12 mph
Downstrm: b + c = 18 mph
------------------------------
Add and solve for "b":
2b = 30 mph
b = 15 mph (boat speed in still water)
---
Solve for "c":
b + c = 18 mph
15 + c = 18
c = 3 mph (speed of the current)
=====================================
Cheers,
Stan H.
==================